OpenStudy (anonymous):

How can I simplify this expression using only cos and sin? cos2(2x)-sin(2x)

OpenStudy (anonymous):

do you mean that??. \[\cos ^{2}(2x)-\sin(2x)\]

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

\[\cos(2x)=2\cos ^{2}x-1\] \[\sin(2x)=2sinx*cosx\] can you continue??

OpenStudy (anonymous):

well I would need to square the first term, but after that, I don't know what the next step wuld be...

OpenStudy (anonymous):

4cos^4(x)-4cos^2(x)=1-2sinxcosx